OpenStudy (anonymous):

Sherry missed the lesson on normal distribution and needs to do her homework. Explain to Sherry how to use the mean and standard deviation of a normal distribution to determine the top 7% of the population.

OpenStudy (anonymous):

@delila4675

OpenStudy (anonymous):

okay here!!!

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

how do I find the mean and standard deviation of 7%?

OpenStudy (anonymous):

ok first of all do u know what standard deviation means?

OpenStudy (anonymous):

when reading questions all ways look for the key words that stand out.

OpenStudy (anonymous):

so do know what standard deviation means?

OpenStudy (anonymous):

uhh not really but isn't it the amount that doesn't really match the rest of the data?

OpenStudy (anonymous):

okay here is the definition In statistics, the standard deviation (SD) (represented by the Greek letter sigma, σ) is a measure that is used to quantify the amount of variation or dispersion of a set of data values.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now ok The question says Sherry missed the lesson on normal distribution and needs to do her homework. Explain to Sherry how to use the mean and standard deviation of a normal distribution to determine the top 7% of the population. We want Φ(x−μσ)=1−0.07=0.93

OpenStudy (anonymous):

thats the standard deviation equation?

OpenStudy (anonymous):

We'd solve for x. Where as μ is the man, and σ is standard deviation.

OpenStudy (anonymous):

Actually "determine the top 7% of the population" is sort of vague. I interpreted it to mean the score at which you would be at the border. The question is not very precise.

OpenStudy (anonymous):

ok do you want to move onto the next one?

OpenStudy (anonymous):

wait no not done.

OpenStudy (anonymous):

First, you use z=x−μσ to find the z score. μ is mean and σ is standard deviation. Next, you would look up z in your z table. P=Φ(z) In this case P is what you find in the table, and it is the probability that your score is under x.

OpenStudy (anonymous):

With the current question, we have P is already provided. When looking at the top 7%, this is the opposite of being in the bottom 93%. So we let P=0.93. We do the z table look up, but instead of looking up z, we find look to see what z value will give us 0.93.

OpenStudy (anonymous):

then we can finally solve for x.

OpenStudy (anonymous):

1.5 i think

OpenStudy (anonymous):

How, try to find a z that will result you in getting 0.97

OpenStudy (anonymous):

I mean 0.9300

OpenStudy (anonymous):

have caulator. and brb.

OpenStudy (anonymous):

ok im so lost

OpenStudy (anonymous):

okay. here http://openstudy.com/study#/updates/5301ad54e4b024fe26e22d29

OpenStudy (anonymous):

that might help.

OpenStudy (anonymous):

brb have to eat but I will be back for next question ok?

OpenStudy (anonymous):

by.

OpenStudy (anonymous):

ok i will tag you in the next one.